OpenStudy (anonymous):

Eliminate the parameter. x = square root of t, y = 4t + 1

OpenStudy (anonymous):

help me please

OpenStudy (anonymous):

will give medal

OpenStudy (anonymous):

\[ x=\sqrt t\implies t=x^2 \] Can you finish it now?

OpenStudy (anonymous):

\[ y=4 t+ 1\\ y=4 x^2 +1 \]

OpenStudy (anonymous):

thanks eli